OpenStudy (anonymous):

Write an equation in slope-intercept form for each line described. 1.) passes through (-12, 3), perpendicular to y= -2/3x-11

OpenStudy (tkhunny):

You should practice this until you can just write it down about as fast as you can read it. Point-Slope (since that what we are given) is (y-y0) = m(x-x0) We know a point, (-12,3). This leads immediately to: (y-3) = m(x-(-12)) We ALMOST know a slope. That annoying "perpendicular" hint gives it to us. If lines are perpendicular and neither is vertical, we can multiply their slopes and get -1. Thus, (-2/3)*m = -1 and m = 3/2. This leads immediately to (y-3) = (3/2)(x-(-12)) and we are almost done. You will have to change it to slope-intercept as requested in the problem statement.

OpenStudy (anonymous):

Perpendicular means m1*m2 = -1 ie the product of slopes = 1